# Bayesian optimisation of poloidal field coil positions in tokamaks

### v1

@@ 3.2.3 Constraints
In addition to the objectives, we also have a number of design and engineering constraints that need to be satisfied so that each PF coil set considered in the BO loop produces an equilibrium with key targets that are similar to the baseline equilibrium and does not violate coil current limits. The bounds of the constraints are summarised in Table 3.1.
@@
The first constraint is on the shape of the LCFS, which is defined as the contour of $(R,Z)$ points that pass through the X-point closest to the magnetic axis—recall Figure 3.1. We denote this region of the $RZ$-plane as $\Omega_{p}$ and quantify the difference between two different regions using
@@
where ||$\cdot$|| ⋅ | denotes the cross-sectional area of a region in the poloidal plane (Bardsley et al., 2024). This parameter quantifies the ratio of the total non-overlapping areas and the sum of the two areas. Placing an upper limit on this ratio enables us to constrain the LCFS shape of the new equilibrium ($\Omega^{2}_{p}$) to be similar to that of the baseline ($\Omega^{1}_{p}$). This helps to ensure the new equilibrium has similar core performance to the baseline.
The second and third constraints place an upper limit on the distance between the strikepoints (i.e. where the separatrix first intersects the limiter geometry at some location [^7] In rare cases, an equilibrium may have a separatrix that intersects the limiter multiple times on the same plate.) and the centre of the strike plates. The bound is half of the length of the strike plate, with one constraint on each of the inner and outer strike plates.
The fourth constraint will place an upper limit on the distance between the two X-point locations when mirrored about $Z=0$. This distance should be minimal in a double-null plasma scenario as considered here. See Figure 3.2 for the strikepoint and X-point locations.
@@
Constraint Bound [unit]
LCFS area ratio $\leqslant 0.012$
Outer strike distance $\leqslant 0.14$ [m]
Inner strike distance $\leqslant 0.32$ [m]
X-point distance $\leqslant 0.01$ [m]
Maximum current density $\leqslant 100$ [$\mathrm{MA/m^{2}}$]
Table 3.1: Constraint bounds enforced on the equilibria generated by FreeGS for a particular PF coil set.
3.2.4 Equilibrium simulator
In order to calculate the aforementioned objective functions and evaluate whether or not the constraints have been met, we need a simulator that is able to generate a plasma equilibrium using the STEP baseline design and a given PF coil set. For this we use FreeGS, a free-boundary static inverse equilibrium solver developed by Dudson (2023). FreeGS will return a plasma equilibrium (in terms of the poloidal flux) and the PF coil currents required to generate it. It uses an optimisation routine to identify the coil currents with respect to some constraints on the chosen plasma shape and a Picard iteration scheme to solve the free-boundary Grad-Shafranov problem (see Song et al. (2024) and Pentland et al. (2024)). The required inputs to solve the equilibrium problem are:
@@ 3.2.4 Equilibrium simulator
  • 23 isoflux constraints [^8] These are constraints that ensure the poloidal flux $\psi(R,Z)$ at two different locations $(R_{1},Z_{1})$ and $(R_{2},Z_{2})$ are the same. that link poloidal flux values on the core plasma boundary to the X-points and the divertor regions.
@@
3.2.5 Failure regions
Figure 3.3: Two examples of invalid equilibria returned by “converged” FreeGS simulations with different PF coil sets.
@@ 3.2.5 Failure regions
In Figure 3.3, we illustrate two cases of invalid equilibria returned by FreeGS. The left panel depicts an equilibrium in which both X-points have formed inside the divertor regions, far from the desired locations in the core, resulting in the LCFS intersecting the limiter. This is problematic when calculating the ICL, as this calculation assumes the LCFS does not intersect the limiter geometry until hitting the strikeplate, resulting in an unfeasibly small ICL value. The right panel shows a single-null equilibrium with the LCFS, again, intersecting the limiter in the core region. Calculations for both the ICL/OCL and the strike distances are spurious in this case.
To mitigate these issues, we can classify (recall Section 2.3) whether an equilibrium is valid by checking the following conditions:
@@
4 Numerical experiments
In this section, we will perform the design optimisation of the PF coil set problem. The first experiment will use multi-objective BO to find several Pareto optimal PF coil sets that respect the engineering and design constraints in Section 3.2. We analyse two of the Pareto optimal in more detail, highlighting how the BO explores the solution space while respecting the trade-off between the objective functions. To further illustrate the data efficiency of the BO, we compare its performance against two other optimisation methods (simple Sobol sampling and a genetic algorithm) when using both identical and larger computational budgets.
To generate these results, we use the Trieste platform (Picheny et al., 2023; Berkeley et al., 2024) which provides the software implementations for Sobol sampling, acquisition functions, and Gaussian processes (via GPflow (Matthews et al., 2017)). Pygmo2 (Biscani and Izzo, 2020) provides the genetic algorithm which we will use for benchmarking. When evaluating the Sobol samples with FreeGS, we make use of the CSD3 HPC cluster (see Acknowledgements) and the Simvue platform (Lahiff et al., 2024) to monitor simulation progress and store the objective/constraint data.
4.1 Stand-alone BO
@@
In Table 4.1, we display the proportions of each sampling set that result in feasible, infeasible (violating one or more constraints), and failed (invalid) PF coil sets. We can see that only $10\%$ of the Sobol samples provide feasible designs and that once the BO loop begins running, we accumulate a much larger proportion of feasible designs (with fewer failures). This shows that the GPs can accurately model the constraint responses and the acquisition function uses this to propose feasible samples. The drop in failure region samples likely results from a combination of explicit classifier intervention and the scarcity of optimal samples near these regions, making them less likely to be chosen by the acquisition function.
@@
Table 4.1: The number (and percentage) of samples from each sampling method that lie in failure, infeasible, or feasible regions.
@@
Figure 4.1: Feasible solutions in the objective space for the $64$ Sobol + $64$ BO experiment. Shown are the Sobol solutions (blue) and the BO solutions (light pink to red). Pareto optimal solutions are denoted with a star and dominated solutions with a circle, with light pink to red indicating successive BO iterations.
Figure 4.2: The total number of feasible BO samples at the end of each BO iteration.
@@
The BO loop successfully identifies the Pareto front (see Figure 4.1), capturing the trade-off between the volumetric sum (3.1) and the average connection length (recall (3.2)). Generally, the later BO iterations produce samples that dominate earlier samples, highlighting how BO learns from new data and exploits its new understanding of the functions to produce higher-quality samples.
@@
During the initial BO iterations, the data is sparse, resulting in highly uncertain and inaccurate GPs. As a result, the exact location of the feasible regions is unclear and, as it explores for the first $20$ iterations, the BO produces few feasible samples—see Figure 4.2. Following this initial exploration, $30$ iterations of exploitation takes place, where BO reliably produces feasible samples (those seen in Figure 4.1). The final $20$ iterations yield no feasible samples, indicating a return to exploration and suggesting there is little scope to find feasible optimal solutions around the current Pareto optimal points.
Objective/constraint Sample $\bm{110}$ Sample $\bm{71}$
$V$ $77.6$ [$\mathrm{m}^{3}$] $85.1$ [$\mathrm{m}^{3}$]
ICL $96.9$ [$\mathrm{m}$] $108$ [$\mathrm{m}$]
OCL $70.3$ [$\mathrm{m}$] $71.8$ [$\mathrm{m}$]
LCFS area ratio $0.0117$ $0.0103$
Outer strike distance $0.0957$ [$\mathrm{m}$] $0.0962$ [$\mathrm{m}$]
Inner strike distance $0.0233$ [$\mathrm{m}$] $0.0305$ [$\mathrm{m}$]
X-point distance $0.00394$ [$\mathrm{m}$] $0.000919$ [$\mathrm{m}$]
$J_{\text{max}}$ $82.4$ [$\mathrm{MA}/\mathrm{m}^{2}$] $98.8$ [$\mathrm{MA}/\mathrm{m}^{2}$]
Inner $\langle B_{p}\rangle$ $0.413$ [$\mathrm{T}$] $0.373$ [$\mathrm{T}$]
Outer $\langle B_{p}\rangle$ $0.537$ [$\mathrm{T}$] $0.530$ [$\mathrm{T}$]
@@
Table 4.2: The objectives and constraint values (to three significant figures) returned for two of the Pareto optimal solutions shown in Figure 4.3. Also shown are the inner and outer line-averaged poloidal magnetic field readings.
Figure 4.3: Two Pareto optimal PF coil sets (and the separatrix of the equilibria they generate) returned from the $71$st (green) and $110$th (green) sampling of the simulator (corresponding to samples from the $7$th and $46$th BO iteration, respectively).
@@
In Table 4.2, we display the objective/constraint values obtained from two of the Pareto optimal PF coil sets shown in Figure 4.3, with both samples obtained during the BO iterations. The $71$st sample yields the the highest average connection length while the $110$th sample yields the lowest volumetric sum (i.e. they are the two samples that optimise the marginals of the objective-space). The first three rows of the table show the objective quantities for both samples while the intermediate five rows show the constraint values. The constraints show that samples $110$ and $71$ are close to the constraint bounds for the LCFS shape and maximum coil current density, respectively. This could indicate that further optimisation of these samples (and other Pareto optimal samples) is not possible without violating the constraints, hence the lack of feasible samples in the final BO iterations above.
@@
It is clear from Figure 4.3 that the $110$th sample has a smaller volumetric sum because PF coils 2, 3, 4, and 5 are closer to the device centreline ($R=0$). However, the difference in average connection length is less obvious because the separatrices look (qualitatively, at least) very similar. The lower connection length in sample $110$ results from a higher poloidal field, causing particles travelling from the midplane into the divertors to move faster, decreasing the number of times (and thus the distance) they orbit the tokamak toroidally. This can be seen in the final two rows of Figure 4.3, which shows the line-averaged poloidal magnetic field $\langle B_{p}\rangle$ along the inner and outer connection length field lines, respectively.
4.2 Comparison between BO, Sobol sampling, and a genetic algorithm
@@
Next, we compare the BO scheme against two other multi-objective optimisation methods. The first method we compare against uses quasi-Monte Carlo sampling via the Sobol method, essentially relying on random chance to sample feasible and optimal coil sets. The second method will use a genetic algorithm, specifically the ‘Non-Dominated Sorting Genetic Algorithm II’ (NSGA-II) (Deb et al., 2002). NSGA-II, like all genetic algorithms, operates on the principle that combining the inputs of well-performing individuals within a population can produce offspring (new samples) that inherit characteristics from its parents and therefore may perform similarly or better. Starting with an initial population (in our case, Sobol samples), the algorithm generates additional samples through iterative recombination and mutation (applying slight random changes to the inputs), therefore introducing variation to explore the solution space (Verma et al., 2021). Constraints are handled by penalising the objectives according to the number of violated constraints (Kuri-Morales and Gutiérrez-García, 2002). Similarly, the failure region is handled by returning large constants for the objectives, artificially making the sample appear very non-optimal.
@@
We run two sets of experiments with each of these methods, the results of which are presented in Table 4.3. The first two (II and III) use the same computational budget as the BO experiment (I) from Section 4.1, while the second two (IV and V) have a budget 8×8$\times$8 × larger.
@@
From the results, we can see that experiment I produces better samples than both II and III with a hypervolume at least $20\%$ larger. Recall that a larger hypervolume indicates a feasible objective space with better trade-offs that are further from the (anti-optimal) reference point—the feasible solutions are shown in the objective space in Figure 4.4. The genetic algorithm finds the most feasible samples, outperforming BO by nearly 2×2$\times$2 ×. However, the hypervolume of experiment III indicates few of these feasible samples offer any improvement over even quasi-random samples. This illustrates how the BO performs significantly better than the Sobol sampling and the genetic algorithm at finding Pareto optimal PF coil sets. Genetic algorithms find the most feasible PF coil sets, however, all of the samples are of significantly lower quality than those from BO.
@@
III 8 Sobol + 120 NSGA-II $1659.2$ $49\%$
IV 1024 Sobol $1711.5$ $13\%$
V 32 Sobol + 992 NSGA-II $1820.3$ $64\%$
Table 4.3: The hypervolume of the feasible solution set and the percentage of total samples taken that were feasible for each of the experiments run. Here, we compare the $64$ Sobol + $64$ BO experiment from Section 4.1 with pure Sobol sampling and the NSGA-II algorithm, each with the same number of samples (128). We also display Sobol and NSGA-II experiments that use 8×8$\times$8 × the number of samples (1024). All hypervolumes are calculated with respect to the same reference point.
Figure 4.4: Feasible solutions in the objective space for experiments I (red), II (blue), and III (green) in Table 4.3. Pareto optimal solutions are denoted with a star and dominated solutions with a circle.
Figure 4.5: Feasible solutions in the objective space for experiments I (red), IV (blue), and V (green) in Table 4.3. Pareto optimal solutions are denoted with a star and dominated solutions with a circle.
BO continues to outperform Sobol sampling and the genetic algorithm even when we increase their computational budgets to $1024$ samples. While the hypervolumes returned in experiments IV and V are larger compared to those in II and III (as expected), they still cannot reach the level achieved by the BO (with $1/8$th of the data). In Figure 4.5, we again see the majority of Pareto optimal samples coming from the BO with a few being found by the genetic algorithm, with BO finding the best samples in each objective (the samples that optimise the marginals of the objective space). The vast majority of samples taken by the alternative methods are, however, dominated by others from the BO. Again, genetic algorithms find the most feasible samples, however, they form a front that underperforms that of BO, particularly in the volumetric sum.
5 Discussion and outlook

### v2

@@ III.2.3 Constraints
In addition to the objectives, we also have a number of design and engineering constraints that need to be satisfied so that each PF coilset considered in the BO loop produces an equilibrium with key targets that are similar to the baseline equilibrium and does not violate coil current limits. The bounds of the constraints are summarised in Table 1.
Constraint Bound [unit]
LCFS area ratio $\leq 0.012$
Outer strike distance $\leq 0.14$ [m]
Inner strike distance $\leq 0.32$ [m]
X-point distance $\leq 0.01$ [m]
Maximum current density $\leq 100$ [$\mathrm{MA/m^{2}}$]
Table 1: Constraint bounds enforced on the equilibria generated by FreeGS for a particular PF coilset.
@@
The first constraint is on the shape of the LCFS, which is defined as the contour of $(R,Z)$ points that pass through the X-point closest to the magnetic axis—recall Figure 2. We denote this region of the $RZ$-plane as $\Omega_{p}$ and quantify the difference between two different regions using
@@
where ||$\cdot$|| ⋅ | denotes the cross-sectional area of a region in the poloidal plane (Bardsley, Baker, and Vincent, 2024). This parameter quantifies the ratio of the total non-overlapping areas and the sum of the two areas. Placing an upper limit on this ratio enables us to constrain the LCFS shape of the new equilibrium ($\Omega^{2}_{p}$) to be similar to that of the baseline ($\Omega^{1}_{p}$). This helps to ensure the new equilibrium has similar core performance to the baseline.
The second and third constraints place an upper limit on the distance between the strikepoints (i.e. where the separatrix first intersects the limiter geometry at some location) and the centre of the strike plates. In rare cases, an equilibrium may have a separatrix that intersects the limiter multiple times on the same plate, hence we need to account for that. The bound is half of the length of the strike plate, with one constraint on each of the inner and outer strike plates.
The fourth constraint will place an upper limit on the distance between the two X-point locations when mirrored about $Z=0$. This distance should be minimal in a double-null plasma scenario as considered here. See Figure 3 for the strikepoint and X-point locations.
@@
III.2.4 Equilibrium simulator
In order to calculate the aforementioned objective functions and evaluate whether or not the constraints have been met, we need a simulator that is able to generate a plasma equilibrium using the STEP baseline design and a given PF coilset. For this we use FreeGS, a free-boundary static inverse equilibrium solverDudson (2023). FreeGS will return a plasma equilibrium (in terms of the poloidal flux) and the PF coil currents required to generate it. It uses an optimisation routine to identify the coil currents with respect to some constraints on the chosen plasma shape and a Picard iteration scheme to solve the free-boundary Grad-Shafranov problem (see Song et al. (2024) and Pentland et al. (2024)). The required inputs to solve the equilibrium problem are:
@@ III.2.4 Equilibrium simulator
  • 23 isoflux constraints that link poloidal flux values on the core plasma boundary to the X-points and the divertor regions (i.e. constraints that ensure the poloidal flux $\psi(R,Z)$ at two different locations $(R_{1},Z_{1})$ and $(R_{2},Z_{2})$ are the same).
@@
III.2.5 Failure regions
Figure 4: Two examples of invalid equilibria returned by “converged” FreeGS simulations with different PF coilsets.
@@ III.2.5 Failure regions
In Figure 4, we illustrate two cases of invalid equilibria returned by FreeGS. The left panel shows a single-null equilibrium with the LCFS intersecting the limiter in the core region. The right panel depicts an equilibrium in which both X-points have formed inside the divertor regions, far from the desired locations in the core, resulting in the LCFS again intersecting the limiter. This is problematic when calculating the ICL, as this calculation assumes the LCFS does not intersect the limiter geometry until hitting the strikeplate, resulting in an unfeasibly small ICL value. Calculations for both the ICL/OCL and the strike distances are spurious in this case.
To mitigate these issues, we can classify (recall Section II.3) whether an equilibrium is valid by checking the following conditions:
@@
IV Numerical experiments
In this section, we will perform the design optimisation of the PF coil set problem. The first experiment will use multi-objective BO to find several Pareto optimal PF coilsets that respect the engineering and design constraints in Section III.2. We analyse two of the Pareto optimal solutions in more detail, highlighting how the BO explores the solution space while respecting the trade-off between the objective functions. To further illustrate the data efficiency of the BO, we compare its performance against two other optimisation methods (simple Sobol sampling and a genetic algorithm) when using both identical and larger computational budgets.
To generate these results, we use the Trieste (Picheny et al., 2023; Berkeley et al., 2024) package which provides the software implementations for Sobol sampling, acquisition functions, and Gaussian processes (via GPflow (Matthews et al., 2017)). Pygmo2 (Biscani and Izzo, 2020) provides the genetic algorithm which we will use for benchmarking. When evaluating the Sobol samples with FreeGS(Dudson, 2023), we make use of the CSD3 HPC cluster (see Acknowledgements) and the Simvue platform (Lahiff et al., 2024) to monitor simulation progress and store the objective/constraint data.
IV.1 Stand-alone BO
@@
In Table 2, we display the proportions of each sampling set that result in feasible, infeasible (violating one or more constraints), and failed (invalid) PF coilsets. We can see that only $10\%$ of the Sobol samples provide feasible designs and that once the BO loop begins running, we accumulate a much larger proportion of feasible designs with fewer failures. This shows that the GPs can accurately model the constraint responses and the acquisition function uses this to propose feasible samples.
The drop in failure region samples likely results from a combination of explicit classifier intervention and the scarcity of optimal samples near these regions, making them less likely to be chosen by the acquisition function. The classifier has a precision of $0.82$, meaning $82\%$ of the area included in the acquisition maximisation (by zeroing the ECHVI in these regions) is indeed non-failing, reducing waste of computational resources by potentially sampling failing points. A recall of $0.86$ shows that only $14\%$ of the non-failure region is incorrectly avoided by the classifier; it is more important that this percentage is low because Pareto optimal solutions could exist here but would be missed.
@@
Table 2: The number (and percentage) of samples from each sampling method that lie in failure, infeasible, or feasible regions.
@@
Figure 5: Feasible solutions in the objective space for the $64$ Sobol + $64$ BO experiment. Shown are the baseline (grey cross), Sobol solutions (blue) and the BO solutions (light pink to red). Pareto optimal solutions are denoted with a star and dominated solutions with a circle, with light pink to red indicating successive BO iterations. Two of the BO samples, $71$ and $110$, are highlighted for further analysis.
Figure 6: The total number of feasible BO samples at the end of each BO iteration.
@@
The BO loop successfully identifies the Pareto front (see Figure 5), capturing the trade-off between the volumetric sum (8) and the average connection length (recall (9)). Generally, the latter BO iterations produce samples that dominate earlier samples, highlighting how BO learns from new data and exploits its new understanding of the functions to produce higher-quality samples.
@@
During the initial BO iterations, the data is sparse, resulting in highly uncertain and inaccurate GPs. As a result, the exact location of the feasible regions is unclear and, as it explores for the first $20$ iterations, the BO produces few feasible samples—see Figure 6. Following this initial exploration, $30$ iterations of exploitation takes place, where BO reliably produces feasible samples (those seen in Figure 5). The final $20$ iterations yield no feasible samples, indicating a return to exploration and suggesting there is little scope to find feasible optimal solutions around the current Pareto optimal points.
Objective/constraint Baseline Sample $\bm{110}$ Sample $\bm{71}$
$V$ [$\mathrm{m}^{3}$] $86.5$ $77.6$ $85.1$
ICL [$\mathrm{m}$] $96.5$ $96.9$ $108$
OCL [$\mathrm{m}$] $67.1$ $70.3$ $71.8$
LCFS area ratio $0.0108$ $0.0117$ $0.0103$
Outer strike distance [$\mathrm{m}$] $0.0934$ $0.0957$ $0.0962$
Inner strike distance [$\mathrm{m}$] $0.0297$ $0.0233$ $0.0305$
X-point distance [$\mathrm{m}$] $0.000374$ $0.00394$ $0.000919$
$J_{\text{max}}$ [$\mathrm{MA}/\mathrm{m}^{2}$] $71.4$ $82.4$ $98.8$
Inner $\langle B_{p}\rangle$ [$\mathrm{T}$] $0.414$ $0.413$ $0.373$
Outer $\langle B_{p}\rangle$ [$\mathrm{T}$] $0.552$ $0.537$ $0.530$
@@
Table 3: The objectives and constraint values (to three significant figures) for the baseline and two of the Pareto optimal solutions shown in Figure 7. Also shown are the inner and outer line-averaged poloidal magnetic field readings.
@@
In Table 3, we display the objective/constraint values obtained from the baseline and two of the Pareto optimal PF coil sets shown in Figure 7, with both samples obtained during the BO iterations. The $71$st sample yields the highest average connection length while the $110$th sample has the joint lowest volumetric sum (of these tied samples, it has the higher connection length). The first three rows of the table show the objective quantities for both samples while the intermediate five rows show the constraint values.
The objective values of the baseline (and its location in Figure 5) relative to the Pareto optimal samples demonstrate that BO is able to improve the PF coilset design significantly. At a minimum, BO has yielded a reduction of $V$ by $1.4\mathrm{m}^{3}$ and an increase in average connection length of $1.8\mathrm{m}$ over the baseline; this is not insignificant considering the low computational budget to achieve these gains. The constraints show that samples $110$ and $71$ are close to the constraint bounds for the LCFS shape and maximum coil current density, respectively. This could indicate that further optimisation of these samples (and other Pareto optimal samples) is not possible without violating the constraints, hence the lack of feasible samples in the final BO iterations.
@@
It is clear from Figure 7 that the $110$th sample has a smaller volumetric sum because PF coils 2, 3, 4, and 5 are closer to the device centreline ($R=0$). However, the difference in average connection length is less obvious because the separatrices look (qualitatively, at least) very similar. The lower connection length in sample $110$ results from a higher poloidal field, causing particles travelling from the midplane into the divertors to move faster, decreasing the number of times (and thus the distance) they orbit the tokamak toroidally. This can be seen in the final two rows of Figure 7, which shows the line-averaged poloidal magnetic field $\langle B_{p}\rangle$ along the inner and outer connection length field lines, respectively.
Figure 7: The PF coilsets and corresponding separatrices for the baseline (pink) and two of the Pareto optimal equilibria: sample $71$ (blue) and $110$ (green) of the simulator (corresponding to samples from the $7$th and $46$th BO iteration, respectively).
IV.2 Comparison between BO, Sobol sampling, and a genetic algorithm
@@
Next, we compare the BO scheme against two other multi-objective optimisation methods. The first method we compare against uses quasi-Monte Carlo sampling via the Sobol method, essentially relying on random chance to sample feasible and optimal coilsets. The second method will use a genetic algorithm, specifically the ‘Non-Dominated Sorting Genetic Algorithm II’ (NSGA-II) (Deb et al., 2002). NSGA-II, like all genetic algorithms, operates on the principle that combining the inputs of well-performing individuals within a population can produce offspring (new samples) that inherit characteristics from its parents and therefore may perform similarly or better. Starting with an initial population (in our case, Sobol samples), the algorithm generates additional samples through iterative recombination and mutation (applying slight random changes to the inputs), therefore introducing variation to explore the solution space (Verma, Pant, and Snasel, 2021). Constraints are handled by penalising the objectives according to the number of violated constraints (Kuri-Morales and Gutiérrez-García, 2002). Similarly, the failure region is handled by returning large constants for the objectives, artificially making the sample appear very non-optimal.
@@
We run six additional experiments with each of these methods, the results of which are presented in Table 4. The first three (II, III and IV) use the same computational budget as the BO experiment (I) from Section IV.1, while the final three (V, VI and VII) have a budget that is 8×8$\times$8 × larger.
Four of the experiments contain results for the NSGA-II genetic algorithm. For each budget, we include an NSGA-II experiment with the same initial population ($64$ samples) as Experiment I (experiments III and VI) and another with a lower initial population (experiments IV and VII) but with the ability to run over more generations. In both cases, the experiments using a lower initial population size (IV and VII) outperform their counterparts with larger initial population sizes. This is because they are able to explore the solution space more widely (using more iterations), thus generating more feasible solutions and making incremental progress each generation towards the Pareto frontier. This can be seen by the higher percentage of feasible solutions sampled by these experiments compared to the others. For the remaining analysis, we will compare only these best performing NSGA-II experiments (IV and VII) against the Sobol sampling and BO.
@@
From the results, we can see that experiment I produces better samples than both II and IV with a hypervolume at least $20\%$ larger. Recall that a larger hypervolume indicates a feasible objective space with better trade-offs that are further from the (anti-optimal) reference point—the feasible solutions are shown in the objective space in Figure 8. The genetic algorithm finds the most feasible samples, outperforming BO by nearly 2×2$\times$2 ×. However, the hypervolume of experiment IV indicates few of these feasible samples offer any improvement over even quasi-random samples. This illustrates how the BO performs significantly better than the Sobol sampling and the genetic algorithm at finding Pareto optimal PF coilsets. Genetic algorithms find the most feasible PF coilsets, however, all of the samples are of significantly lower quality than those from BO.
@@
III 64 Sobol + 64 NSGA-II $1544.2$ $16\%$
IV 8 Sobol + 120 NSGA-II $1659.2$ $49\%$
V 1024 Sobol $1711.5$ $13\%$
VI 64 Sobol + 960 NSGA-II $1713.6$ $23\%$
VII 32 Sobol + 992 NSGA-II $1820.3$ $64\%$
Table 4: The hypervolume of the feasible solution set and the percentage of total samples taken that were feasible for each of the experiments run. Here, we compare the $64$ Sobol + $64$ BO experiment from Section IV.1 with pure Sobol sampling and the NSGA-II algorithm, each with the same number of samples (128). We also display Sobol and NSGA-II experiments that use 8×8$\times$8 × the number of samples (1024). All hypervolumes are calculated with respect to the same reference point.
Figure 8: Feasible solutions in the objective space for experiments I (red), II (blue), and IV (green) in Table 4. Pareto optimal solutions are denoted with a star and dominated solutions with a circle.
Figure 9: Feasible solutions in the objective space for experiments I (red), V (blue), and VII (green) in Table 4. Pareto optimal solutions are denoted with a star and dominated solutions with a circle.
BO continues to outperform Sobol sampling and the genetic algorithm even when we increase their computational budgets to $1024$ samples. While the hypervolume returned in experiments V and VII are larger compared to those in II and IV (as expected), they still cannot reach the level achieved by the BO (with $1/8$th of the data). In Figure 9, we again see the majority of Pareto optimal samples coming from the BO with a few being found by the genetic algorithm, with BO finding the best samples in each objective (the samples that optimise the marginals of the objective space). The vast majority of samples taken by the alternative methods are, however, dominated by others from the BO. Again, genetic algorithms find the most feasible samples, however, they form a front that underperforms that of BO, particularly in the volumetric sum.
V Discussion and outlook